Barb is buying a new car for $14,000. Her old car has a trade-in Value of $2500. The dealer informs her that the financing charge is 3% add-on interest. She wishes to take 2.5 years to pay off the car. Answer the following:

How much is her down payment? [trade-in value of old car]

How much does she need to finance?

How much interest will she pay?

What will her monthly payment be?

What will be the total amount to be repaid?
Answer #1

ANSWER:

WORKING
A DOWN PAYMENT (TRADE IN VALUE OF OLD CAR) $2,500
B AMOUNT SHE NEEDS TO FINANCE =14000-2500 $11,500
C TOTAL INTEREST SHE WILL PAY IN 2.5 YEARS =11500*0.03*30/12 862.5
ASSUMING INTEREST IS PER ANNUM BASIS
D MONTHLY PAYMENT PRINCIPAL =11500/30 $383.33
INTEREST =H6/30 $28.75
TOTAL MONTHLY PAYMENT $412.08
E TOTAL AMOUNT TO BE REPAID PRINCIPAL =14000-2500 $11,500
INTEREST =11500*0.03*30/12 862.5
TOTAL AMOUNT $12,363